# WaveStitch: Flexible and Fast Conditional Time Series Generation with Diffusion Models

### v2

@@ Abstract.
Generating temporal data under conditions is crucial for forecasting, imputation, and generative tasks. Such data often has metadata and partially observed signals that jointly influence the generated values. However, existing methods face three key limitations: (1) they condition on either the metadata or observed values, but rarely both together; (2) they adopt either training-time approaches that fail to generalize to unseen scenarios, or inference-time approaches that ignore metadata; and (3) they suffer from trade-offs between generation speed and temporal coherence across time windows—choosing either slow but coherent autoregressive methods or fast but incoherent parallel ones. We propose WaveStitch, a novel diffusion-based method to overcome these hurdles through: (1) dual-sourced conditioning on both metadata and partially observed signals; (2) a hybrid training-inference architecture, incorporating metadata during training and observations at inference via gradient-based guidance; and (3) a novel pipeline-style paradigm that generates time windows in parallel while preserving coherence through an inference-time conditional loss and a stitching mechanism. Across diverse datasets, WaveStitch demonstrates adaptability to arbitrary patterns of observed signals, achieving 1.81x lower mean-squared-error compared to the state-of-the-art, and generates data up to 166.48x faster than autoregressive methods while maintaining coherence. Our code is available at: https://github.com/adis98/WaveStitch
@@
submissionid: 280copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: ACM Management of Data; May 31–June 05, 2026; Bangalore, Indiaccs: Information systems Data management systems
@@ 1. Introduction
Figure 1. Yearly revenue (signal) grouped by metadata (District, Brand, Month, etc.). We want to support different conditional scenarios, like C1: (District A + Brand Y), or C2: (District B + Month August). Observed signals are in green.
Conditional synthesis allows injecting user-defined constraints into the generation process, such as provided metadata or partially observed signals (Narasimhan et al., 2024; Kollovieh et al., 2024; Alcaraz and Strodthoff, 2023). For instance, we can use it for data management tasks such as imputation (Chen et al., 2023), forecasting (Faloutsos et al., 2018; Qiu et al., 2024), and data cleaning (Chu et al., 2016) by providing clean parts of the data as conditional observations for the model to regenerate or correct the rest. We can also augment data or simulate new scenarios for downstream tasks (Liu et al., 2024; Sanghi and Haritsa, 2023; Narasimhan et al., 2024). While existing works view forecasting as a special case of imputation, where the missing values occur only at the tail-end of the time series  (Alcaraz and Strodthoff, 2023; Kollovieh et al., 2024), generative tasks are more loosely specified. Instead of matching exact values, generative tasks aim to produce plausible and diverse samples that are consistent with high-level constraints such as metadata.
@@
Challenge 1. Existing methods fail to condition on both the metadata (e.g., brand, district) and the observed signals (shown in green in Figure 1) in conjunction. Figure 1 shows how data can exhibit varying patterns, such as contiguous missing blocks (C1), or fine-grained gaps (C2). Existing models focus either on the observed values and neglect the metadata (Alcaraz and Strodthoff, 2023; Kollovieh et al., 2024), or vice versa (Narasimhan et al., 2024; Suh et al., 2024). Conditioning only on metadata can capture global trends, but would fail to align with the observed signals. Conversely, conditioning only on the observed signals may locally fill in gaps but ignore the broader context from the metadata. The design challenge lies in combining these heterogeneous conditioning sources—discrete metadata and real-valued observations—while accounting for their differing roles in shaping global and local temporal patterns.
Challenge 2. The second challenge is deciding when to incorporate conditional information—during training or inference. The choice is influenced by the differing characteristics of metadata and observed values. Metadata are static, hence the model can learn their conditional influence during training. In contrast, signals may be observed at arbitrary timesteps (see C1 and C2 in Figure 1), requiring flexibility at inference. However, existing works adopt a black-or-white stance, conditioning either during training (Narasimhan et al., 2024; Tashiro et al., 2021; Rasul et al., 2021; Suh et al., 2024; Alcaraz and Strodthoff, 2023) or during inference (Kollovieh et al., 2024; Lugmayr et al., 2022).
Challenge 3. Time series can be generated either autoregressively or in parallel. Autoregression incurs high runtime costs, generating signals sequentially using sliding windows to propagate temporal dependencies across time windows (Rasul et al., 2021) (see Figure 2). In contrast, parallel methods generate multiple windows simultaneously (Kollovieh et al., 2024; Narasimhan et al., 2024), but ignore temporal coherence across windows. The challenge lies in maintaining efficiency and coherence—a balance that current methods struggle to achieve.
@@
Our key novelty is a unified framework to handle all three challenges. We dual-source our conditioning strategy to use metadata and observed signals. This combination requires hybridising training and inference. We first train a base generative model that conditions only on metadata. The observed signals are injected directly at inference, guiding the sampling trajectory using gradients from a conditional loss. This loss incorporates a novel stitching mechanism to enforce coherence across overlapping time-windows, restricting the generation process to remain close to the realistic space. The approach is similar to gradient inversion (Jeon et al., 2021), where gradients update the sample to match a given objective while keeping the model frozen. Moreover, the conditional loss is easily parallelised, enabling fast synthesis. This innovation allows us to break away from prior approaches and introduce a new paradigm for time series generation through pipeline parallelism. Multiple time windows are generated simultaneously with dependencies propagating across overlaps, maintaining high efficiency and temporal coherence.
@@
1. Dual-sourced Conditioning: WaveStitch combines two complementary sources of information: metadata that captures global context (Sec. 3.3), and observed signals (Sec. 3.4) that provide fine-grained local cues to refine generation with high precision.
2. Hybrid Architecture: WaveStitch combines the strengths of training and inference-time strategies. It conditions on metadata during training (Sec. 3.3), and incorporates the observed values at inference via gradient-based guidance using a conditional loss (Sec. 3.4). This enables generalisation to arbitrarily positioned patterns while preserving the global structure dictated by metadata.
3. Coherent Pipelined-Parallel Generation: To improve efficiency, we generate consecutive time windows in a pipelined-parallel fashion, incorporating a novel stitching mechanism within the conditional loss to reconcile overlaps at inference. As windows are denoised simultaneously, dependency information propagates across overlaps along the pipeline, ensuring smooth and coherent transitions without compromising speed (Secs. 3.44).
4. Evaluations on Diverse Tasks: We evaluate WaveStitch on point-wise and synthetic data quality metrics, demonstrating its flexibility across tasks, achieving up to 1.81x lower Mean-Squared-Error (MSE) on average compared to the state-of-the-art (SOTA) baselines. Results show that WaveStitch ’s parallel method is significantly faster (166.48x) than autoregressive synthesis while maintaining generation quality. Additional ablations assess the impact of the conditioning strategy, stitch loss formulations, and random imputation tasks with varying degrees of missingness (Sec. 5).
@@ 2. Background and Problem Definition
We cover foundations for diffusion models for time series and their conditional extensions, and formally define the problem while briefly highlighting the research gaps. Table 1 summarises all the key symbols and notations used henceforth.
@@ 2.1. Diffusion Models for Time Series
Denoising Diffusion Probabilistic Models (DDPMs) are a class of generative models used extensively for time series (Ho et al., 2020; Alcaraz and Strodthoff, 2023; Kollovieh et al., 2024; Yuan and Qiao, 2024). As shown in Figure 3, DDPMs operate by first gradually corrupting the data with a noising process known as forward diffusion—and then learn to reverse this through a backward denoising process composed of stochastic transitions (Ho et al., 2020). We forward noise a time window of $w$ timesteps starting from timestep $i$, as follows [^1] There are two types of ”steps” defined in this study. Diffusion steps refer to the iterative noise addition or removal in the generative process, while timesteps refer to the temporal axis in the time series data.:
@@
Figure 3. Forward-Backward diffusion. $\mathbf{x}^{(:)}_{0}$ and $\mathbf{x}^{(:)}_{T}$ are the initial signal window and fully noised window after $T$ steps
where $\mathbf{x}^{(i:i+w-1)}_{t}$ is the multivariate time series signal after noising $t\in\{0,1\dots T\}$ steps, $1-\alpha_{t}$ is the $t$-th noise variance, (i:i+w1)(i:i+w-1) indicates the time slice222both $i$ and i+w1i+w-1 are included, yielding window size $w$, and $\epsilon_{t}\sim\mathcal{N}(0,I)$ is sampled white noise. Given this notation, $\mathbf{x}^{(:)}_{0}$, $\mathbf{x}^{(:)}_{T}$ represent a clean and fully noised sample for an arbitrary window slice, and $T$ is the number of diffusion steps. Noising up to any step without recursively applying Eq. (1) is done with a reparameterization trick (Ho et al., 2020):
@@
where $\hat{\mathbf{x}}^{(:)}_{t}$ and $\hat{\mathbf{x}}^{(:)}_{t-1}$ represent the signal estimates at step $t$ and t1t-1 respectively. $\sigma_{t}$ is a function of $\alpha$ to ensure stochasticity during sampling, and $z\sim\mathcal{N}(0,I)$ (Ho et al., 2020). We train the denoiser by noising the signal up to a random step $t$ using Eq. (2), and then use the MSE loss for backpropagation, similar to (Ho et al., 2020). The aim is to push the denoised sample towards matching the ground truth, i.e., $\hat{x}_{0}^{(:)}\simeq x_{0}^{(:)}$.
@@ 2.2. Conditional Diffusion for Time Series
Standard DDPMs cannot generate samples adhering to conditions, as synthetic samples are generated freely from random noise. In contrast, conditional DDPMs constrain the denoising process—using metadata or observed signals (Alcaraz and Strodthoff, 2023; Narasimhan et al., 2024; Kollovieh et al., 2024). The conditions can be incorporated at different times—during training or inference.
@@ 2.2.1. Training-time Conditioning
There are three main ways to condition a DDPM during training. The first way is to provide the conditions as additional inputs when training the model—either in the form of metadata (e.g. TimeWeaver (Narasimhan et al., 2024)), or binary masks indicating the missing and conditional observations, like SSSD  (Alcaraz and Strodthoff, 2023) and CSDI (Tashiro et al., 2021). The second way is classifier-free guidance (Ho and Salimans, 2021; Suh et al., 2024), which trains both a conditional and an unconditional denoiser, and interpolates between their outputs to control the degree of conditioning. The third way is classifier guidance (Liu et al., 2024; Dhariwal and Nichol, 2021), where an external classifier or controller guides the sampling of an unconditional denoiser(Liu et al., 2024). The classifier acts like a feedback loop to adjust the denoiser’s trajectory by estimating the likelihood of meeting the conditions given the current output state.
@@ 2.2.2. Inference-time Conditioning
These approaches control sampling directly at inference by correcting or ”guiding” an unconditional diffusion model’s outputs. RePaint (Lugmayr et al., 2022; Chen et al., 2023), adds a controlled amount of noise to the observed signals, to resemble the noise-corrupted samples seen during training. An unconditional model then denoises the entire sequence iteratively, generating both the observed and missing values jointly. Self-Guidance  (Kollovieh et al., 2024) conditions on arbitrary patterns of observed values by leveraging the unconditional denoiser’s own estimates. Unlike classifier guidance, which relies on a separately trained classifier to estimate likelihoods, self-guidance repurposes the denoiser itself for this task. It first computes a rough signal estimate $\hat{\mathbf{x}}^{(:)}_{0}$ from the current noisy state $\hat{\mathbf{x}}^{(:)}_{t}$, by inverting the forward noising process in Eq. (2). The samples are then iteratively adjusted using the gradient of the conditional log-likelihood of the observed signals directly at inference.
(a) Root Level Condition: (B,,)(B,*,*)
(b) Intermediate Condition: (,January,)(*,{January},*)
(c) Bottom Condition: (,,15)(*,*,15)
Figure 4. Coarse-grained (a), Intermediate (b), and Fine-grained (c) tasks in Example 1.
@@ 2.3. Problem Definition
We are given a time series dataset $\mathcal{D}=\{(\mathbf{a}^{(i)},\mathbf{x}^{(i)})\}_{i=1}^{N},$ consisting of $N$ timesteps, where at the $i$th timestep, $\mathbf{a}^{(i)}\in\mathbb{R}^{1\times L}$ denotes a sample with $L$ metadata features and $\mathbf{x}^{(i)}\in\mathbb{R}^{1\times C}$ is the $C$-variate signal. The aim is to generate time series signals given the metadata and arbitrary patterns of signal observations.
Training inputs include complete pairs of metadata and signals. That is, training is done on signals $\mathbf{x}^{(i)}$ whose metadata $\mathbf{a}^{(i)}$ are known. The model must generate time series signals conditioned on the metadata and the observed signals at inference time. Conditional metadata may be specified as follows: $\mathbf{a}^{(k)}=(\text{val}_{1},\text{val}_{2},*,*,*)$, indicating that signals should be generated (or possibly re-generated) at all timesteps $k$ where the first two metadata columns match $\text{val}_{1}$ and $\text{val}_{2}$. The asterisk (*) indicates unconstrained features. Therefore, the set $\{\mathbf{a}^{(k)}\}$ is assumed to contain all valid enumerations of the unspecified metadata features that satisfy the given condition. These enumerations are then inserted into $\mathcal{D}$ at their appropriate positions in the timeline. The goal is to then generate signals $\mathbf{x}^{(k)}$ conditioned on $\mathbf{a}^{(k)}$ and any observed signals $\mathbf{x}^{(i\neq k)}$ from $\mathcal{D}$.
We further illustrate the problem of conditioning on metadata and observed signals in Example 1 and Figure 4.
@@
[innerleftmargin = 5pt] Example 1. Consider a dataset containing one year’s worth of daily sales data for three brands—A, B, and C. Each Sales entry (the signal) corresponds to a particular Brand, on a given Month, and Day. The data is organised by concatenating the entries from all three brands into a table where each row is assigned a timestep index based on its position.
A condition at the root metadata feature, such as $(\text{\emph{Brand}}=B,*,*)$ requires generating sales for the entire year for Brand B. The timesteps needing generation are shown in 4(a). The generated portions comprise a long contiguous block of timesteps, i.e., coarse-grained granularity. In contrast, a condition on the bottom metadata feature, such as $(*,*,\text{\emph{Day}}=15)$ generates sales for the $15^{\text{th}}$ day of each month and brand. This scenario resembles fine-grained tasks with the generated values dispersed across timesteps, as shown in 4(c). Tasks can also have intermediate granularity, like 4(b) .
Table 1. Summary of Key Symbols
@@
$\mathbf{x}^{(:)}_{t},\mathbf{\hat{x}}^{(:)}_{t}$ Signal and denoised windows at $t$th diffusion step
@@
$N\text{ or }M,L,C,T$ #Timesteps, #metadata, #channels, #diffusion steps
@@
$\mathbf{x}^{(i)},\mathbf{m}^{(i)},\mathbf{a}^{(i)}$ Signal, mask, and metadata at $i$th timestep
$\mathbf{x}_{w}^{(j)},\mathbf{m}_{w}^{(j)},\mathbf{a}_{w}^{(j)}$ $j$th signal, mask, and metadata window samples
$\mathbf{\hat{x}}_{w,t}^{(j)}$ $j$th signal window at diffusion step $t$

### v3

@@ Abstract.
Generating temporal data under conditions is crucial for forecasting, imputation, and generative tasks. Such data often has metadata and partially observed signals that jointly influence the generated values. However, existing methods face three key limitations: (1) they condition on either the metadata or observed values, but rarely both together; (2) they adopt either training-time approaches that fail to generalise to unseen scenarios, or inference-time approaches that ignore metadata; and (3) they suffer from trade-offs between generation speed and temporal coherence across time windows, choosing either slow but coherent autoregressive methods or fast but incoherent parallel ones. We propose WaveStitch, a novel diffusion-based method to overcome these hurdles through: (1) dual-sourced conditioning on both metadata and partially observed signals; (2) a hybrid training-inference architecture, incorporating metadata during training and observations at inference via gradient-based guidance; and (3) a novel pipeline-style paradigm that generates time windows in parallel while preserving coherence through an inference-time conditional loss and a stitching mechanism. Across diverse datasets, WaveStitch demonstrates adaptability to arbitrary patterns of observed signals, achieving 1.81x lower mean-squared-error compared to the state-of-the-art, and generates data up to 166.48x faster than autoregressive methods while maintaining coherence. Our code is available at: https://github.com/adis98/WaveStitch.
@@
submissionid: 280copyright: ccjournal: PACMMODjournalyear: 2025journalvolume: 3journalnumber: 6 (SIGMOD)article: 377publicationmonth: 12doi: 10.1145/3769842ccs: Information systems Data management systems
@@ 1. Introduction
Figure 1. Time series (signal) of partially observed yearly revenues (in green), grouped by metadata District, Brand and Month. We want to generate missing data under varying conditions, like C1: (District A + Brand Y), or C2: (District B + Month August).
Conditional synthesis enables injecting user-defined constraints into the generation process, such as provided metadata or partially observed signals (Narasimhan et al., 2024; Kollovieh et al., 2024; Alcaraz and Strodthoff, 2023). For example, we can use it for data management tasks such as imputation (Chen et al., 2023), forecasting (Faloutsos et al., 2018; Qiu et al., 2024), and data cleaning (Chu et al., 2016) by providing clean parts of the data as conditional observations for the model to regenerate or correct the rest. For such tasks, forecasting is often treated as a special case of imputation, where missing values occur only at the tail-end of the time series  (Alcaraz and Strodthoff, 2023; Kollovieh et al., 2024). Going beyond imputation and forecasting, we can also have generative tasks that are more loosely defined and are used for tasks like data augmentation or simulating new scenarios (Liu et al., 2024; Sanghi and Haritsa, 2023; Narasimhan et al., 2024). Unlike forecasting and imputation, generative tasks target producing samples that are consistent and plausible with high-level constraints such as metadata, rather than having an exact match with partially observed signals.
@@
Challenge 1. Existing methods fail to condition on both the metadata (e.g., brand, district) and the observed signals in conjunction. Figure 1 shows that data can exhibit varying missingness patterns, such as contiguous blocks (C1), or fine-grained gaps (C2). Existing models condition only on the observed values and ignore the metadata (Alcaraz and Strodthoff, 2023; Kollovieh et al., 2024), or vice versa (Narasimhan et al., 2024). Using only metadata can capture global trends, but would fail to align with the observed signals. Conversely, conditioning only on the observed signals may locally fill in gaps but overlook the broader context given by the metadata. The design challenge lies in using both sources to account for their differing influence in shaping global and local temporal patterns.
Challenge 2. The second challenge is deciding when to incorporate conditional information, during training or inference. The choice is influenced by the differing characteristics of metadata and observed values. Metadata is static, and a model can learn its conditional influence during training. In contrast, signals may be observed at arbitrary timesteps (see C1 and C2 in Figure 1), requiring flexibility at inference. However, existing works adopt a black-or-white stance, conditioning exclusively during training (Narasimhan et al., 2024; Tashiro et al., 2021; Rasul et al., 2021; Suh et al., 2024; Alcaraz and Strodthoff, 2023) or during inference (Kollovieh et al., 2024; Lugmayr et al., 2022).
Challenge 3. Time series signals can be generated either autoregressively or in parallel. Autoregression incurs high runtime costs, generating signals sequentially using a sliding window to propagate temporal dependencies across consecutive segments (Rasul et al., 2021) (see Figure 2). In contrast, parallel methods generate multiple windows simultaneously (Kollovieh et al., 2024; Narasimhan et al., 2024), but ignore maintaining temporal coherence between them. The challenge lies in maintaining both efficiency and coherence, a balance that current methods struggle to achieve.
@@
Our key novelty is a unified framework to handle all three challenges. WaveStitch dual-sources its conditioning to use both metadata and observed signals. This combination requires hybridising training and inference. We first train a base generative model that conditions only on metadata. We inject the observed signals directly at inference, guiding the sampling trajectory using gradients from a conditional loss. This loss incorporates a novel stitching mechanism to enforce coherence across overlapping time windows, thereby restricting the generation process to remain close to the realistic space. The approach is similar to gradient inversion (Jeon et al., 2021), where gradients update the sample to match a given objective while keeping the model frozen. Moreover, the conditional loss is easily parallelised, enabling fast synthesis. This innovation allows us to break away from prior approaches and introduce a new paradigm for time series generation through pipeline parallelism. Multiple time windows are generated simultaneously with dependencies propagating across overlaps, maintaining high efficiency and temporal coherence.
@@
1. Dual-sourced Conditioning: WaveStitch combines two complementary sources of information: metadata for capturing global context (Sec. 3.3), and observed signals (Sec. 3.4) for providing fine-grained local cues to refine generation with high precision.
2. Hybrid Architecture: WaveStitch combines the strengths of training and inference-time strategies. It learns the conditional influence of the (static) metadata during training (Sec. 3.3), and dynamically refines samples at inference using any observed signal values via gradient-based guidance (Sec. 3.4). This approach enables the handling of arbitrary patterns of observed signals while preserving the global structure dictated by the metadata.
3. Coherent Pipelined-Parallel Generation: To improve efficiency, we generate consecutive time windows in a pipelined-parallel fashion, incorporating a novel stitching mechanism to reconcile overlaps at inference. Dependency information propagates across time windows while they are generated in parallel. This process resembles a pipeline, ensuring smooth and coherent transitions without compromising speed (Secs. 3.44).
4. Evaluations on Diverse Tasks: We evaluate WaveStitch on point-wise and synthetic data quality metrics, demonstrating its flexibility across tasks and achieving up to 1.81x lower Mean-Squared-Error (MSE) compared to the state-of-the-art (SOTA) baselines. Results show that WaveStitch’s parallelism enables significantly faster generation than autoregressive synthesis (166.48x) while maintaining generation quality. Additional ablations assess the impact of the conditioning strategy, stitch loss formulations, and random imputation tasks with varying degrees of missingness (Sec. 5).
@@ 2. Background and Problem Definition
We cover the foundations of diffusion models for time series, their conditional extensions, and formally define the problem while briefly highlighting the research gaps. Table 1 summarises all the key symbols and notations used henceforth.
@@ 2.1. Diffusion Models for Time Series
Denoising Diffusion Probabilistic Models (DDPMs) are a class of generative models used extensively for time series (Alcaraz and Strodthoff, 2023; Kollovieh et al., 2024; Yuan and Qiao, 2024; Narasimhan et al., 2024). As shown in Figure 3, DDPMs operate by gradually corrupting data with a forward (noising) process, and then learning to reverse this through a backward denoising process (Ho et al., 2020). We forward noise a time window of $w$ timesteps starting from timestep $i$, as follows [^1] There are two types of ”steps” defined in this study. Diffusion steps refer to the iterative noise addition or removal in the generative process, while timesteps refer to the temporal axis in the time series data.:
@@
Figure 3. Diffusion process. $\mathbf{x}^{(:)}_{0}$ and $\mathbf{x}^{(:)}_{T}$ represent the clean and fully noised time window respectively.
where $\mathbf{x}^{(i:i+w-1)}_{t}$ is the multivariate time series signal after noising $t\in\{0,1\dots T\}$ steps, $1-\alpha_{t}$ is the $t$-th noise variance, (i:i+w1)(i:i+w-1) indicates the time slice222both $i$ and i+w1i+w-1 are included, yielding window size $w$, and $\epsilon_{t}\sim\mathcal{N}(0,I)$ is sampled white noise. With this notation, $\mathbf{x}^{(:)}_{0}$ and $\mathbf{x}^{(:)}_{T}$ represent a clean and fully noised sample, respectively. Noising up to any step without recursively applying eq. (1) is done with a reparameterisation trick (Ho et al., 2020):
@@
where $\hat{\mathbf{x}}^{(:)}_{t}$ and $\hat{\mathbf{x}}^{(:)}_{t-1}$ represent the signal estimates at step $t$ and t1t-1 respectively, $\sigma_{t}$ is a function of $\alpha$ to ensure sample diversity, and $z\sim\mathcal{N}(0,I)$ (Ho et al., 2020). The denoiser trains by noising the signal up to a random step $t$ using eq. (2), and then uses the MSE loss with the sampled noise for backpropagation, pushing the sample towards matching the ground truth, i.e., $\hat{x}_{0}^{(:)}\simeq x_{0}^{(:)}$ (Ho et al., 2020).
@@ 2.2. Conditional Diffusion for Time Series
Standard DDPMs cannot generate samples adhering to conditions, as synthetic samples are generated freely from random noise in an unconstrained manner. In contrast, conditional DDPMs constrain the denoising process using metadata or observed signals (Alcaraz and Strodthoff, 2023; Narasimhan et al., 2024; Kollovieh et al., 2024). These conditions can be incorporated at different times, either during training or inference.
@@ 2.2.1. Training-time Conditioning
There are three main ways to condition a DDPM during training. The first way is to provide the conditions as additional inputs (e.g. metadata (Narasimhan et al., 2024)) during training, or through binary masks to indicate the timesteps with missing and observed signals  (Alcaraz and Strodthoff, 2023; Tashiro et al., 2021). The second way is classifier-free guidance (Ho and Salimans, 2021; Suh et al., 2024), which trains both a conditional and an unconditional model, and interpolates between their outputs to control the degree of conditioning. The third way is classifier guidance (Liu et al., 2024; Dhariwal and Nichol, 2021), where an external classifier or controller guides the sampling of an unconditional denoiser (Liu et al., 2024). The classifier functions as a feedback loop, adjusting the model’s outputs by estimating the likelihood of meeting the given conditions at each denoising step.
@@ 2.2.2. Inference-Time Conditioning
These approaches control sampling directly at inference by steering or guiding the outputs of an unconditional diffusion model. RePaint (Lugmayr et al., 2022; Chen et al., 2023), adds a controlled amount of noise to the observed signals, to resemble the noise-corrupted samples seen during training. An unconditional model then denoises the entire sequence, generating both the observed and missing values jointly. Self-Guidance  (Kollovieh et al., 2024) conditions on arbitrary patterns of observed values by leveraging the unconditional model’s own estimates. Unlike classifier guidance, which relies on a separately trained classifier for likelihood estimation, self-guidance repurposes the denoiser itself for this task. It first computes a rough signal estimate $\hat{\mathbf{x}}^{(:)}_{0}$ from the current outputs $\hat{\mathbf{x}}^{(:)}_{t}$, by inverting the forward noising process in eq. (2). The samples are then iteratively corrected using the gradient of the conditional log-likelihood of the observed signals directly at inference.
(a) Root: (B,,)(B,*,*)
(b) Intermediate: (,January,)(*,{January},*)
(c) Bottom: (,,15)(*,*,15)
Figure 4. (a) Root (coarse-grained), (b) Intermediate and (c) Bottom level (fine-grained) conditions in Example 1.
@@ 2.3. Problem Definition
We are given a time series dataset. $\mathcal{D}=\{(\mathbf{a}^{(i)},\mathbf{x}^{(i)})\}_{i=1}^{N}$ with $N$ timesteps, where $\mathbf{a}^{(i)}\in\mathbb{R}^{1\times L}$ denotes a sample with $L$ metadata features and $\mathbf{x}^{(i)}\in\mathbb{R}^{1\times C}$ is a $C$-variate signal at the $i^{\text{th}}$ timestep. The aim is to generate time series signals given the metadata and signals observed at arbitrary timesteps. Training inputs include complete pairs of metadata and signals, i.e., training uses signals $\mathbf{x}^{(i)}$ with available metadata $\mathbf{a}^{(i)}$. The model must generate time series signals conditioned on the metadata and the observed signals at inference time. Conditional metadata may be specified as $\mathbf{a}^{(k)}=(\text{u},\text{v},*,*,*)$, indicating that signals should be generated (or possibly re-generated) at all timesteps $k$ where the first two metadata columns match u and v respectively. The asterisk (*) indicates unconstrained features. Therefore, the set $\{\mathbf{a}^{(k)}\}$ is assumed to contain all valid enumerations of the unspecified metadata features that satisfy the given condition. These enumerations are then inserted into $\mathcal{D}$ at their appropriate positions in the timeline. The model then needs to generate signals $\mathbf{x}^{(k)}$ conditioned on $\mathbf{a}^{(k)}$ and any observed signals $\mathbf{x}^{(i\neq k)}$ already present in $\mathcal{D}$. We further illustrate this problem setup using Example 1 and Figure 4.
@@
[innerleftmargin = 5pt] Example 1. Consider a dataset containing one year’s worth of daily sales data for three brands, A, B, and C. Each Sales entry (the signal) corresponds to a particular Brand, on a given Month, and Day. The data is organised by concatenating the entries from all three brands into a table where each row is assigned a timestep index based on its position.
A condition on the root-level metadata feature, such as $(\text{\emph{Brand}}=B,*,*)$, requires generating sales for the entire year for Brand B. As 4(a) shows, the generated portions comprise a long contiguous block of timesteps, resembling coarse-grained synthesis. In contrast, a condition on the bottom-level metadata feature, such as $(*,*,\text{\emph{Day}}=15)$, generates sales for the $15^{\text{th}}$ day of each month and brand, resembling fine-grained tasks, as shown in 4(c). Similarly, tasks may also have intermediate granularity by conditioning mid-level features, like in 4(b) .
Table 1. Summary of key symbols.
@@
$\mathbf{x}^{(:)}_{t},\mathbf{\hat{x}}^{(:)}_{t}$ Ground-truth signal and model estimate at the $t$th diffusion step
@@
N $\text{ or }M,L,C,T N$umber of timesteps, metadata features, signal channels, and diffusion steps
@@
$\mathbf{x}^{(i)},\mathbf{m}^{(i)},\mathbf{a}^{(i)}$ Signal (ground-truth), mask, and metadata at the $i$th timestep
$\mathbf{x}_{w}^{(j)},\mathbf{m}_{w}^{(j)},\mathbf{a}_{w}^{(j)}$ $j$th signal window (ground-truth), mask, and metadata window
$\mathbf{\hat{x}}_{w,t}^{(j)}$ $j$th signal window (model estimate) at diffusion step $t$